OpenStudy (anonymous):

im trying to find side length b of a right triangle, where a is 23, and c is 83. i think you need the pathagorean therom or w.e it is. but i cant seem to get the right answer. is it 106, 86.13 79.75 or 60?

OpenStudy (anonymous):

what r a and c?

OpenStudy (anonymous):

mean u must know what is base, perpendicular and hyp

OpenStudy (anonymous):

a is 23 and c is 83. so i square a and square c then subtract the a squared from c squared to get b squared then get the square root of that but i think thats wrong because it doesnt work.

OpenStudy (anonymous):

If I use your calculation then b = 79.75 (I assume c = hypotenuse).

OpenStudy (anonymous):

a^2+b^2=c^2 so we have 23^2+b^2=83^2 b^2=83^2-23^2 = \[2*\sqrt{1590} = 79.74961 which rounds \to 79.75\]